# Dataset Distillation-based Hybrid Federated Learning on Non-IID Data

### v2

@@ V-C1 Server
In HFLDD, the server is responsible for client clustering and global model aggregation. In the process of client clustering, the server needs to compute the similarity matrix $M$, which involves the computation of KL Divergence between any pair of clients’ soft labels. As shown in Eq. 4, for each pair of clients $i$ and $j$, the complexity of computing $M_{ij}$ over the global dataset $D_{g}$ is $\mathcal{O}(|D_{g}|\cdot N_{c})$. In total, the complexity of constructing the similarity matrix $M$ is $\mathcal{O}(N^{2}\cdot|D_{g}|\cdot N_{c})$. Based on the similarity matrix $M$, the server conducts homogeneous client clustering using K-Means algorithm, which incurs a complexity of $\mathcal{O}(I\cdot K\cdot N^{2})$, where $I$ is the number of iterations in K-means. In the process of global model aggregation, the server aggregates models from all $H$ cluster heads over $T$ communication rounds. Each aggregation involves weighted averaging $H$ local models with size being $|\omega|$, resulting in a total computational cost of $\mathcal{O}(T\cdot H\cdot|\omega|)$.
@@ V-C2 Cluster Member
Each cluster member performs pre-training and soft label generation at the stage of label knowledge collection and dataset distillation at the stage of IID dataset generation. In the process of local pre-training, each client performs $E_{1}$ steps of model updates on its local dataset using mini-batch SGD. Let $b_{1}$ denote the mini-batch size used in each step. The computational complexity of local pre-training per client is $\mathcal{O}(E_{1}\cdot b_{1}\cdot|\omega|)$. After pre-training, each client generates a soft label tensor by performing inference over the public dataset $\mathcal{D}_{g}$, which incurs an additional computational cost of $\mathcal{O}(|\mathcal{D}_{g}|\cdot|\omega|)$. Therefore, the computation cost per client at the label knowledge collection stage is $\mathcal{O}(E_{1}\cdot b_{1}\cdot|\omega|+|\mathcal{D}_{g}|\cdot|\omega|)$. In the process of dataset distillation, each cluster member synthesizes distilled data via the KIP method, which involves repeated optimization over $R$ iterations. In each iteration, for cluster member $i$, $i\in\mathcal{H}_{e,h}$ and $i\notin\mathcal{A}$, the most computationally intensive step is the inversion of the regularized kernel matrix in kernel ridge regression, which incurs a cost of $\mathcal{O}(|\tilde{\mathcal{D}_{i}}|^{3})$. Therefore, the complexity of dataset distillation at cluster member $i$ is $\mathcal{O}(R\cdot|\tilde{\mathcal{D}_{i}}|^{3}).$
@@ V-C3 Cluster Head
In HFLDD, at the stage of label knowledge collection, each cluster head follows the same pre-training and soft label generation steps as cluster members, which incurs a computational complexity of $\mathcal{O}(E_{1}\cdot b_{1}\cdot|\omega|+|\mathcal{D}_{g}|\cdot|\omega|)$. At the stage of global model training, each cluster head additionally conducts local training on its hybrid dataset. In each communication round, the cluster head at cluster $h$ performs $E_{2}$ steps of mini-batch SGD on its dataset $\mathcal{D}_{h}$. Let $b_{2}$ denote the batch size. The computational complexity of local training per round on a single cluster head is $\mathcal{O}(E_{2}\cdot b_{2}\cdot|\omega|)$. For all $T$ rounds, the total complexity is $\mathcal{O}(T\cdot E_{2}\cdot b_{2}\cdot|\omega|)$.
@@ VI-A1 Datasets
Our experiments are carried out on three widely used datasets MNIST [49], CIFAR10 [50], and CINIC10 [51]. MNIST consists of 50,000 binary images of handwritten digits, which can be divided into 10 classes. CIFAR10 consists of 50,000 daily object images, which can also be divided into 10 classes. CINIC10, as an extended version of CIFAR10, comprises 270,000 images and retains the original 10 classes of CIFAR10. By incorporating downsampled images from ImageNet, CINIC10 significantly increases the complexity and diversity of the dataset. In our non-IID setting, we investigate scenarios in which each client has 1, 2, or 10 of the ten classes. In our experiments, each dataset is divided into 80% for training and 20% for testing.
In addition to training and testing the local client dataset, we also need additional global datasets to generate soft labels for label knowledge collection. We randomly select 1,000 samples from other similar datasets and construct the corresponding global dataset. For the MNIST experiment, the global dataset is constructed from the FMNIST dataset [52]. For the CIFAR10 and CINIC10 experiments, the global dataset is constructed from the Tinyimagenet dataset [53].
@@ VI-A2 Models
In our experiments, the applied models are all based on Convolutional Neural Network (CNN) [3]. For the MNIST dataset, we use a shallow CNN consisting of two convolutional layers with ReLU activations and max pooling, followed by flattening and three fully connected layers. The model contains 44,426 parameters. For the CIFAR10 dataset, a deeper CNN is adopted with four convolutional blocks. Each block includes two convolutional layers with ReLU activations, batch normalization, and a max pooling layer. After flattening, a single fully connected layer is applied. The total number of parameters is 1,020,160. For the CINIC10 dataset, the model consists of three repeated convolutional blocks, each comprising a convolutional layer with 128 filters, followed by an instance normalization layer, a ReLU activation function, and an average pooling layer. The total number of model parameters is 320,010. The input for all models is $32\times 32\times 3$ image and the output is a 10-dimensional one-hot class label. It should be noted that the proposed HFLDD is not limited by the model architectures, and it can be applied to any model. In the experiments, we also test the performance of HFLDD using different models.
@@ VI-A3 Baselines
  • FedAvg [11], a classic FL framework, where each client conducts local training and the server aggregates all the local models for global model update;
  • FedProx [11], a classic regularization-base FL framework, which adopt regularization terms to adjust local updates, and mitigate client drift on non-IID data;
  • FedSeq [34], a classic HFL framework, which utilizes client clustering and sequential training to enhance model performance;
  • DynaFed [42], a dataset distillation-based FL framework, which synthesizes distilled data by matching the training trajectory of the global model on the server, and then fine-tunes the aggregated model using the generated data
@@ VI-A4 Implementation Details
Our experiments are carried out on a machine equipped with a Tesla V100 GPU and implemented using PyTorch, we fixed the number of clients $N$ to 100. For the pre-training phase, both the MNIST, CIFAR10 and CINIC10 datasets were used with the following settings: learning rate $\eta=0.01$, local update steps in the pre-training phase $E_{1}=10$, a batch size of 64, and the SGD optimizer. These settings ensure that each client can effectively extract knowledge from its local data, thereby enhancing the discriminability for subsequent clustering tasks. To construct the heterogeneous topology, we set the clustering of K-Means with $K=10$. The size of the distilled dataset, denoted by $|\tilde{\mathcal{D}_{i}}|$, is set to 400, which means that each client generates 400 synthetic samples. Additional compression mechanisms are not applied to the distilled data beyond the dataset distillation process itself. The learning rate for optimizing the distilled data is 0.004, and the maximum number of distillation iterations is 3000. In each iteration, a minibatch of 10 samples is drawn from the original dataset to compute the kernel regression loss and update the distilled data. The regularization coefficient for kernel loss is fixed at $\lambda=1\times 10^{-6}$. Following the transmission of the distilled data, the global model training begins. Each cluster head conducts local training with $E_{2}=2$ local update rounds and a batch size of 32. The server performs global model updates for $T=300$ rounds. All other configurations are consistent with the pretraining phase.
To ensure a fair comparison, FedAvg and FedSeq are configured with the same number of global update rounds and local update epochs as HFLDD. Furthermore, in the FedSeq algorithm, the number of clusters is set to 10, with each cluster containing 10 clients. For DynaFed, the server synthesizes 15 distilled data per class. The global model is then fine-tuned for 20 epochs using a batch size of 256. For centralized learning, the training batch size is 64, and the learning rate is 0.01.
@@ VI-B Training Performance
To comprehensively show the training performance of the proposed HFLDD, we compare the HFLDD with the baseline methods in both the non-IID and IID scenarios.
@@
In Fig. 3, we compare the performance of the HFLDD training with the baseline methods on the datasets of MNIST, CIFAR10, and CINIC10, respectively. When $C=1$, as shown in Fig. 3a, Fig. 3d, and Fig. 3g, it can be observed that our proposed HFLDD algorithm outperforms FedAvg, FedSeq, and DynaFed in all datasets. Specifically, on the CIFAR10 dataset, HFLDD achieves an accuracy of 74.0% after 300 communication rounds, significantly outperforming FedAvg at 27.0%, FedProx at 26.3%, FedSeq at 19.6%, and DynaFed at 55.7%. In the CINIC10 dataset, HFLDD achieves an accuracy of 56.7%. Compared to FedAvg, FedProx, FedSeq, and DynaFed, it shows relative improvements of 39.3%, 57.1%, 8.8%, and 41.4%, respectively. On the MNIST dataset, although the final accuracy of HFLDD is similar to that of FedSeq, it is clear that HFLDD converges faster. DynaFed also achieves a similar final accuracy, but the use of distilled data introduces instability in the later stages of training. When $C=2$, on the CIFAR10 dataset, as shown in Figs. 3e and 3g, HFLDD achieves final test accuracies of 78.5% on the CIFAR10 dataset and 59.5% on the CINIC10 dataset, respectively, both of which outperform FedAvg, FedProx, FedSeq and DynaFed. In the MNIST data set, as shown in Fig. 3b, HFLDD, FedSeq, and DynaFed achieve nearly identical training results, and all outperform FedAvg and FedProx. Additionally, under the non-IID settings described above, the performance gap between HFLDD and CL is the smallest.
@@
Figure 5: Performance comparison of HFLDD without dataset distillation (HFLDD w/o DD) with baseline methods on three datasets when $C=10$.
@@ VI-B2 IID Scenario
In this case, we set $C=10$, where each client holds all ten data classes, and the data labels among all clients are balanced. which significantly reduces the benefit of cross-client knowledge aggregation. On the MNIST dataset, as shown in Fig. 3c, all methods achieve nearly identical test accuracies. In the CIFAR10 and CINIC10 dataset, as shown in Fig. 3f and Fig. 3i, HFLDD exhibits slightly inferior performance compared to some baseline methods. However, it remains competitive in the IID setting.
This performance gap can be attributed to two main factors. First, HFLDD trains the global model using a combination of the local data from the cluster head and the distilled data collected from its cluster members. The dataset distillation process inevitably introduces information loss. It is important to emphasize that this information loss is not unique to the IID case but exists across all levels of data heterogeneity. In non-IID settings ($C=1$ or $C=2$), HFLDD still outperforms the baseline methods even with relatively small distilled datasets. To further investigate the impact of dataset distillation, we conduct an experiment in Section VI-D by varying the size of the distilled dataset under different levels of heterogeneity.
Second, under IID condition where each client holds data from all 10 classes, as shown in Figs. 4c, 4f and 4i, the number of clients in each heterogeneous cluster is imbalanced. After cluster heads collect the distilled data, each of them holds data that clearly include all 10 classes, meaning the label distributions across cluster heads are identical. However, due to the different number of clients in each cluster, the amount of data held by each cluster head differs, which leads to quantity skew (each cluster head contains a different number of samples). As discussed in previous work [46], such an imbalance has a minor effect on performance when weighted averaging is used during aggregation. To further assess its influence in our framework, we designed an ablation experiment in which the distilled data transmission step was replaced by direct transmission of the original local data to the cluster head. This variant, denoted by HFLDD w/o DD, removes the influence of dataset distillation. As shown in Fig. 5, HFLDD w/o DD achieves a performance comparable to or even better than the baseline methods in all three datasets, demonstrating that the impact of the cluster size imbalance caused by clustering is indeed minor.
The above analysis and experimental evidence indicate that the observed performance gap between HFLDD and other baseline methods under IID condition is mainly caused by the information loss introduced by dataset distillation, rather than the clustering mechanism.
TABLE II: Communication cost when achieving given accuracy on MNIST under IID condition
@@
TABLE III: Communication cost when achieving given accuracy on CIFAR10 under IID conditions
@@
TABLE IV: Communication cost when achieving given accuracy on CINIC10 under IID conditions
@@
Figure 6: Communication traffic consumption of three methods when achieving given target accuracy on MNIST, CIFAR10 and CINIC10 under non-IID conditions $(C=2)$
@@ VI-C Communication Cost
The analytical communication cost of HFLDD is detailed in Section VI-B. In this part, we compare the communication cost of HFLDD with that of baseline methods, including FedAvg, FedProx, FedSeq and DynaFed.
The communication cost of FedAvg over $T$ rounds of training is given by Eq. (3). The communication costs of FedProx and DynaFed over $T$ rounds are the same as that of FedAvg. For FedSeq, the communication cost comes from parameter sharing between clients within cluster and across clusters, it can be expressed as
$Cost_{FedSeq}\,\,=\,\,O\cdot|\omega|\cdot B_{1}\cdot\left(\left(2T-1\right)+T\cdot J\right),$ (19)
where $O$ is the number of cluster and $J$ is the number of clients in each cluster.
Considering that when the training converges, the communication rounds and test accuracy of different methods are different, it would be unfair to compare the communication cost under a given communication rounds. Therefore, we compare the number of communication rounds and total communication cost required by different methods to achieve a certain test accuracy.
In the IID scenario, the target test accuracies for the MNIST, CIFAR10, and CINIC10 datasets are set to 80%, 70%, and 60%, respectively. Table II to Table IV present the number of communication rounds and total communication cost required by different methods to reach these target accuracies. The ratio indicates the relative communication cost of each method compared to HFLDD. Experimental results show that HFLDD incurs slightly higher communication cost than FedSeq on MNIST and CINIC10, but remains lower than that of FedAvg, FedProx, and DynaFed. On CIFAR10, the communication cost of HFLDD is the lowest among all the methods.
In the Non-IID scenario, when $C=1$, due to the severe imbalance of data labels, the baseline methods may not converge and it is infeasible to find proper target test accuracy that sufficiently reflects the communication cost of each method. Therefore, we consider the case when $C=2$. Fig. 6 compares the communication costs of different methods to achieve given target accuracies on the MNIST, CIFAR10 and CINIC10 datasets. Since the communication cost of HFLDD is significantly lower than that of the baseline methods in this scenario, we use a logarithmic scale (base 10) on the y-axis. From the figure, it can be observed that as the target accuracy increases, HFLDD consistently maintains the lowest communication cost.

### v3

@@ V-C1 Server
In HFLDD, the server is responsible for client clustering and global model aggregation. In the process of client clustering, the server needs to compute the similarity matrix $M$, which involves the computation of KL Divergence between any pair of clients’ soft labels. As shown in Eq. (4), for each pair of clients $i$ and $j$, the complexity of computing $M_{ij}$ over the global dataset $D_{g}$ is $\mathcal{O}(|D_{g}|\cdot N_{c})$. In total, the complexity of constructing the similarity matrix $M$ is $\mathcal{O}(N^{2}\cdot|D_{g}|\cdot N_{c})$. Based on the similarity matrix $M$, the server conducts homogeneous client clustering using K-Means algorithm, which incurs a complexity of $\mathcal{O}(N_{iter}\cdot K\cdot N^{2})$, where $N_{iter}$ is the number of iterations in K-means. In the process of global model aggregation, the server aggregates models from all $H$ cluster heads over $T$ communication rounds. Each aggregation involves weighted averaging $H$ local models with size being $|\omega|$, resulting in a total computational cost of $\mathcal{O}(T\cdot H\cdot|\omega|)$.
@@ V-C2 Cluster Member
Each cluster member performs pre-training and soft labels generation at the stage of label knowledge collection and dataset distillation at the stage of IID dataset generation. In the process of local pre-training, each client performs $E_{1}$ steps of model updates on its local dataset using SGD. Let $b_{1}$ denote the mini-batch size used in each step. The computational complexity of local pre-training per client is $\mathcal{O}(E_{1}\cdot b_{1}\cdot|\omega|)$. After pre-training, each client generates soft labels by performing inference over the public dataset $\mathcal{D}_{g}$, which incurs an additional computational cost of $\mathcal{O}(|\mathcal{D}_{g}|\cdot|\omega|)$. Therefore, the computation cost per client at the label knowledge collection stage is $\mathcal{O}(E_{1}\cdot b_{1}\cdot|\omega|+|\mathcal{D}_{g}|\cdot|\omega|)$. In the process of dataset distillation, each cluster member synthesizes distilled data via the KIP method, which involves repeated optimization over $R$ iterations. In each iteration, for cluster member $i$, $i\in\mathcal{H}_{e,h}$ and $i\notin\mathcal{A}$, the most computationally intensive step is the inversion of the regularized kernel matrix in kernel ridge regression, which incurs a cost of $\mathcal{O}(|\tilde{\mathcal{D}_{i}}|^{3})$. Therefore, the complexity of dataset distillation at cluster member $i$ is $\mathcal{O}(R\cdot|\tilde{\mathcal{D}_{i}}|^{3}).$
@@ V-C3 Cluster Head
At the stage of label knowledge collection, each cluster head follows the same pre-training and soft labels generation steps as cluster members, which incurs a computational complexity of $\mathcal{O}(E_{1}\cdot b_{1}\cdot|\omega|+|\mathcal{D}_{g}|\cdot|\omega|)$. At the stage of global model training, each cluster head additionally conducts local training on its hybrid dataset. In each communication round, the cluster head $a_{h}$ performs $E_{2}$ steps of SGD on its hybrid dataset $\mathcal{D}_{h}$. Let $b_{2}$ denote the batch size. The computational complexity of local training per round on a single cluster head is $\mathcal{O}(E_{2}\cdot b_{2}\cdot|\omega|)$. For all $T$ rounds, the total complexity at the global model training stage is $\mathcal{O}(T\cdot E_{2}\cdot b_{2}\cdot|\omega|)$.
@@ VI-A1 Datasets
Our experiments are carried out on three widely used datasets: MNIST [lecun1998gradient], CIFAR10 [krizhevsky2009learning], and CINIC10 [cinic10]. MNIST consists of 50,000 binary images of handwritten digits, which can be divided into 10 classes. CIFAR10 consists of 50,000 daily object images, which can also be divided into 10 classes. CINIC10, as an extended version of CIFAR10, comprises 270,000 images and retains the original 10 classes of CIFAR10. By incorporating downsampled images from ImageNet, CINIC10 significantly increases the complexity and diversity of the dataset. In our non-IID setting, we investigate scenarios with each client has 1, 2, or 10 of the ten classes. In our experiments, each dataset is divided into 80% for training and 20% for testing.
In addition to training and testing the local client dataset, we also need additional global datasets to generate soft labels for label knowledge collection. We randomly select 1,000 samples from other similar datasets and construct the corresponding global dataset. For the MNIST experiment, the global dataset is constructed from the FMNIST dataset [xiao2017fashion]. For the CIFAR10 and CINIC10 experiments, the global dataset is constructed from the Tinyimagenet dataset [le2015tiny].
@@ VI-A2 Models
In our experiments, the applied models are all based on Convolutional Neural Network (CNN) [krizhevsky2012imagenet]. For the MNIST dataset, we use a shallow CNN consisting of two convolutional layers with ReLU activations and max pooling, followed by flattening and three fully connected layers. The model contains 44,426 parameters. For the CIFAR10 dataset, a deeper CNN is adopted with four convolutional blocks. Each block includes two convolutional layers with ReLU activations, batch normalization, and a max pooling layer. After flattening, a single fully connected layer is applied. The total number of parameters is 1,020,160. For the CINIC10 dataset, the model consists of three repeated convolutional blocks, each comprising a convolutional layer with 128 filters, followed by an instance normalization layer, a ReLU activation function, and an average pooling layer. The total number of model parameters is 320,010. The input for all models is $32\times 32\times 3$ image and the output is a 10-dimensional one-hot class label. It should be noted that the proposed HFLDD is not limited by the model architectures, and it can be applied to any model. In the experiments, we also test the performance of HFLDD using different models.
@@ VI-A3 Baselines
  • FedAvg [mcmahan2017communication], a classic FL framework, where each client conducts local training and the server aggregates all the local models for global model update;
  • FedProx [li2020federated], a classic regularization-base FL framework, which adopt regularization terms to adjust local updates, and mitigate client drift on non-IID data;
  • FedSeq [chen2023fedseq], a classic HFL framework, which utilizes client clustering and sequential training to enhance model performance;
  • DynaFed [pi2023dynafed], a dataset distillation-based FL framework, which synthesizes distilled data by matching the training trajectory of the global model on the server, and then fine-tunes the aggregated model using the generated data
@@ VI-A4 Implementation Details
[…]
To ensure a fair comparison, FedAvg and FedSeq are configured with the same number of global update rounds and local update epochs as HFLDD. Furthermore, in FedSeq, the number of clusters is set to 10, with each cluster containing 10 clients. For DynaFed, the server synthesizes 15 distilled data per class. The global model is then fine-tuned for 20 epochs using a batch size of 256. For centralized learning, the training batch size is 64, and the learning rate is 0.01.
@@ VI-B Training Performance
To comprehensively show the training performance of the proposed HFLDD, we compare HFLDD with the baseline methods in both non-IID and IID scenarios.
@@ VI-B1 Non-IID Scenario
In Fig. 3, we compare the training performance of HFLDD with the baseline methods on the datasets of MNIST, CIFAR10, and CINIC10, respectively. When $C=1$, as shown in Fig. 3a, Fig. 3d, and Fig. 3g, it can be observed that our proposed HFLDD outperforms FedAvg, FedSeq, and DynaFed on all datasets. Specifically, on CIFAR10, HFLDD achieves an accuracy of 74.0% after 300 communication rounds, significantly outperforming FedAvg at 27.0%, FedProx at 26.3%, FedSeq at 19.6%, and DynaFed at 55.7%. On CINIC10, HFLDD achieves an accuracy of 56.7%. Compared to FedAvg, FedProx, FedSeq, and DynaFed, it shows relative improvements of 39.3%, 57.1%, 8.8%, and 41.4%, respectively. On MNIST, although the final accuracy of HFLDD is similar to that of FedSeq, it is clear that HFLDD converges faster. DynaFed also achieves a similar final accuracy, but the use of distilled data introduces instability in the later stages of training. When $C=2$, as shown in Fig. 3e and Fig. 3h, HFLDD achieves final test accuracies of 78.5% on CIFAR10 and 59.5% on CINIC10, respectively, both outperform FedAvg, FedProx, FedSeq and DynaFed. On MNIST, as shown in Fig. 3b, HFLDD, FedSeq, and DynaFed achieve nearly identical training results, and all outperform FedAvg and FedProx. This phenomenon is mainly attributed to the simplicity of the MNIST dataset, which enables different methods to easily learn discriminative features. Additionally, under the non-IID settings described above, the performance gap between HFLDD and CL is the smallest.
@@
Figure 5: Performance comparison of HFLDD without dataset distillation (HFLDD w/o $D$) with baseline methods on three datasets when $C=10$.
@@ VI-B2 IID Scenario
In this case, we set $C=10$, where each client holds all ten data classes, and the data labels among all clients are balanced, which significantly reduces the benefit of cross-client knowledge aggregation. On MNIST, as shown in Fig. 3c, all methods achieve nearly identical test accuracies, which can also be attributed to the simplicity of data. On CIFAR10 and CINIC10, as shown in Fig. 3f and Fig. 3i, HFLDD exhibits slightly inferior performance compared to some baseline methods. However, it remains competitive in the IID setting.
This performance gap can be attributed to two main factors. First, HFLDD trains the global model using a combination of the local data from the cluster head and the distilled data collected from its cluster members. The dataset distillation process inevitably introduces information loss. This effect becomes more noticeable in the IID scenario, where baseline methods can already exploit balanced local data. It is important to emphasize that this information loss is not unique to the IID case but exists across all levels of data heterogeneity. In non-IID settings ($C=1$ or $C=2$), HFLDD still outperforms the baseline methods even with relatively small distilled data. The impact of distilled data size will be shown in Section VI-D.
Second, under IID condition where each client holds data from all 10 classes, as shown in Figs. 4c, 4f and 4i, the number of clients in each heterogeneous cluster is imbalanced. After cluster heads collect the distilled data, each of them clearly holds all 10 classes of data, meaning that the label distributions across cluster heads are identical. However, due to the different number of clients in each cluster, the amount of data held by each cluster head differs, which leads to quantity skew. As discussed in previous work [li2022federated], such an imbalance has a minor effect on performance when weighted averaging is used during aggregation. To further assess its influence in our framework and to isolate the impact of dataset distillation, we conduct an ablation experiment where the distilled data transmission step was replaced by direct transmission of the original local data to the cluster head. This variant, denoted by HFLDD w/o DD, removes the influence of dataset distillation. As shown in Fig. 5, HFLDD w/o DD achieves a performance comparable to or even better than the baseline methods in all three datasets, demonstrating that the impact of the cluster size imbalance caused by clustering is indeed minor.
The above analysis and experimental evidence indicate that the observed performance gap between HFLDD and other baseline methods under the IID condition is primarily attributable to the information loss from dataset distillation, rather than the clustering mechanism.
TABLE III: Communication cost when achieving given accuracy on MNIST under IID condition
@@
TABLE IV: Communication cost when achieving given accuracy on CIFAR10 under IID conditions
@@
TABLE V: Communication cost when achieving given accuracy on CINIC10 under IID conditions
@@
Figure 6: Communication traffic consumption of three methods when achieving given target accuracy under non-IID conditions $(C=2)$.
@@ VI-C Communication Cost
The analytical communication cost of HFLDD is detailed in Eq. (V-B), which accounts for soft labels uploading, distilled data transmission within each cluster, and communication between the server and cluster heads. In this part, we compare the communication cost of HFLDD with that of baseline methods, including FedAvg, FedProx, FedSeq, and DynaFed.
The communication costs of FedProx and DynaFed over $T$ rounds are the same as those of FedAvg, which is given by Eq. (3). For FedSeq, the communication cost comes from parameter sharing between clients within cluster and across clusters, it can be expressed as
$\mathcal{C}_{\textbf{FedSeq}}\,\,=\,\,O\cdot|\omega|\cdot B_{1}\cdot\left(\left(2T-1\right)+T\cdot J\right),$ (19)
where $O$ is the number of clusters and $J$ is the number of clients in each cluster.
Considering that when training converges, the communication rounds and test accuracies of different methods are different, it would be unfair to compare the communication cost under a given communication rounds. Therefore, we compare the number of communication rounds and the total communication cost required by different methods to achieve a certain test accuracy.
In the IID scenario, the target test accuracies for MNIST, CIFAR10, and CINIC10 are set to 80%, 70% and 60%, respectively. Table III to Table V present the number of communication rounds and the total communication costs required by different methods to achieve these target accuracies. The ratio indicates the relative communication cost of each method compared to the HFLDD. Experimental results show that HFLDD incurs slightly higher communication cost than FedSeq on MNIST and CINIC10, but remains lower than that of FedAvg, FedProx, and DynaFed. On CIFAR10, the communication cost of HFLDD is the lowest among all the methods.
In the Non-IID scenario, when $C=1$, due to the severe imbalance of data labels, the baseline methods may not converge and it is infeasible to find a proper target test accuracy that adequately reflects the communication cost of each method. Therefore, we consider the case when $C=2$. Fig. 6 compares the communication costs of different methods to achieve given target accuracies on MNIST, CIFAR10, and CINIC10. Since the communication cost of HFLDD is significantly lower than that of the baseline methods in this scenario, we use a logarithmic scale (base 10) on the y-axis. From the figure, it can be observed that as the target accuracy increases, HFLDD consistently maintains the lowest communication cost.